OpenStudy (anonymous):

If the product of two positive numbers is 1 & their difference is 5/6, then the larger number is.. Help?

OpenStudy (anonymous):

xy = 1 x - y = 5/6 Solve x = 3/2, y = 2/3 The Larger Number is 3/2

OpenStudy (anonymous):

Thank youu! But wait; howd you know it was 3/2 & 2/3?

OpenStudy (anonymous):

\[x(x+\frac{5}{6})=1\]

OpenStudy (anonymous):

\[x^2+\frac{5}{6}x=1\] \[6x^2+5x-6=0\] \[(2x+3)(3x-2)=0\] \[x=-\frac{3}{2},x=\frac{2}{3}\]

OpenStudy (anonymous):

so answers are \[\frac{2}{3}\] and \[\frac{3}{2}\]

OpenStudy (anonymous):

Well that makes alot more sence! Thank youu(: